OpenStudy (anonymous):

Sena chose a number. She multiplied it by 12, added 40, then divided by 8. The result was 26. What was Sena's number?

OpenStudy (campbell_st):

start with 26 then reverse the operations 26 x 8 = then subtract 40 then divide by 12

OpenStudy (johnweldon1993):

Let the number = x \[\frac{ x \times 12 + 40 }{ 8 } = 26\] Can you solve for 'x'...?

OpenStudy (anonymous):

thanks